OpenStudy (anonymous):

What is the slope of the line between (3, –4) and (–2, 1)?

OpenStudy (anonymous):

What is the change in y coordinates??

OpenStudy (anonymous):

3

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

1--4= -3

OpenStudy (anonymous):

See, -(-) = + So, 1 - (-4) becomes 1 + 4 getting??

OpenStudy (anonymous):

5 but how

OpenStudy (anonymous):

-(-) becomes positive... Do you get that??

OpenStudy (anonymous):

ok so how to find the ans

OpenStudy (anonymous):

so you found y change = 5.. Now similarly tell me the change in x coordinates..

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

Yes.. Now, \[Slope = \frac{y \; change}{x \; change}\] Can you do it now??

OpenStudy (anonymous):

0

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

\(Slope = \frac{5}{-5}\) Is it 0???

OpenStudy (anonymous):

no u r wrong

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

Aren't you getting the answer (-1) ???

OpenStudy (anonymous):

is in between -2 and 2

OpenStudy (anonymous):

See, the other formula of slope is: \[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

\[Slope = \frac{1 - (-4)}{-2-3} = \frac{5}{-5} = -1\]

OpenStudy (anonymous):

Is -1 not between 2 and -2??

OpenStudy (anonymous):

\(-2, -1, 0 , 1, 2\) It is between 2 and -2..